# Structurally Flexible Neural Networks: Evolving the Building Blocks for General Agents

### v1

@@ Abstract.
Artificial neural networks used for reinforcement learning are structurally rigid, meaning that each optimized parameter of the network is tied to its specific placement in the network structure. It also means that a network only works with pre-defined and fixed input- and output sizes. This is a consequence of having the number of optimized parameters being directly dependent on the structure of the network. Structural rigidity limits the ability to optimize parameters of policies across multiple environments that do not share input and output spaces. Here, we evolve a set of neurons and plastic synapses each represented by a gated recurrent unit (GRU). During optimization, the parameters of these fundamental units of a neural network are optimized in different random structural configurations. Earlier work has shown that parameter sharing between units is important for making structurally flexible neurons We show that it is possible to optimize a set of distinct neuron- and synapse types allowing for a mitigation of the symmetry dilemma. We demonstrate this by optimizing a single set of neurons and synapses to solve multiple reinforcement learning control tasks simultaneously.
@@ 1. Introduction
The introduction of deep neural networks has led to remarkable improvements when it comes to the control of artificial agents either through optimization using reinforcement learning (Mnih et al., 2013; Arulkumaran et al., 2017) or evolutionary algorithms (Salimans et al., 2017; Stanley et al., 2019). One key finding has been that more diversity in the training environments leads to policies with stronger generalization capabilities (e.g., (Won and Lee, 2019; Kumar et al., 2021; Kirsch et al., 2022b). In the ideal case, we would optimize our agents across a wide array of training environments with differing features, including different input and output spaces. However, most neural networks are structurally rigid. A consequence of this is that their architectures are tied to the specifics of a particular input and output space, which prevents them from being optimized across domains with differing dimensions.
A key challenge on the road to achieving this goal is overcoming what we refer to as the Symmetry Dilemma. The Symmetry Dilemma is the tension between, on one hand, gaining desirable properties that come from inducing symmetry in the optimized parameters, and on the other hand maintaining a structure that is capable of distributing information across its nodes. If we optimize parameters to form a symmetric structure, meaning that any parameter could be placed anywhere within a network structure without changing the functionality of the network, we gain structurally flexible properties such as invariance to permutations and changes in the sizes of the input and/or output (Tang and Ha, 2021; Kirsch et al., 2022a; Kirsch et al., 2022b). However, to distribute information within the structure, each parameter must be able to contain or express individualized information.
@@
In this paper, we present our approach called Structurally Flexible Neural Networks (SFNN). The main features of these networks are that they consist of sparsely connected parameterized neurons and dynamic synapses. The optimized parameters are thus those of the building blocks of a neural network, separated from the network structure. Special for SFNNs, the network contains several different sets of parameterized building blocks. These are referred to as different types of neurons and synapses. The key to achieving a versatile learning algorithm lies in structural flexibility, which liberates the model from being tied to specific input and output spaces. The ultimate objective is to devise a set of building blocks, that when put together to form a network exhibit fast adaptation across various environments, regardless of their dimensionality, and the permutations of the input and output elements. Recently, the notion of foundation models has surfaced in the context of reinforcement learning (Yang et al., 2023). A structurally flexible model could serve as a general foundation model (Bommasani et al., 2021; Yang et al., 2023) for reinforcement learning (RL)-type tasks, even when the tasks do not share input and output dimensions.
In the experiments below, we present promising results showcasing a single parameter set capable of solving tasks with different input and output sizes. The optimized model demonstrates its proficiency in control tasks. Starting with random connectivity, and random initial synaptic weight values, the evolved plastic synapses and neurons organize during the lifetime to achieve better performance in the given environment.
@@ 2. Related Work
Plasticity as Indirect Encoding. Our approach uses Gated Recurrent Units (GRUs) (Chung et al., 2014) with shared parameters to update synaptic weight values. As such, the approach is related to approaches that optimize plasticity mechanisms rather than directly optimizing weights. Indirect encodings offer a powerful way to compress the specification of large systems - like the weights of a neural network - using smaller functions (Zador, 2019; Pedersen and Risi, 2021). From this perspective, approaches that optimize plasticity mechanisms rather than directly optimizing weights can be seen as a way to encode multiple realizations of the parameters of a neural network since these can take different values at different points during the lifetime of an agent (Najarro and Risi, 2020).
@@
Graph Neural Networks. The network type presented in this paper consists of synapses that are themselves recurrent neural networks. This means that they all have hidden state vectors associated with them, and, as explained in Section 3.1, these are updated using local information as well as a global reward signal. This is similar to how Graph Neural Networks (GNNs) work. GNNs are designed to analyze data that is represented as graphs (Wu et al., 2020). Nodes in a GNN have internal states and pass information to their neighbors through parameterized message functions that are optimized during the network’s training phase (Zhou et al., 2020). A neural network with parameterized neurons and synapses being GRUs can be seen as a graph where some nodes (the synapses) can only have two neighbors (neurons), whereas the ”neuron” nodes can have many ”synaptic” nodes as neighbors. In the networks of this paper, these different types of nodes also differ in how they propagate information and update their states. However, the idea of optimizing how nodes with internal states interact with each other is analogous to that of GNNs. GNNs have been used to control simulated robots showing that they can take advantage of the fact that robot morphology can be expressed as a graph (Wang et al., 2018). In contrast, the networks presented in this paper do not make assumptions about the spatial structure of inputs or outputs of the model but are only concerned with how the units of the network can best be updated.
@@ 3. Structurally Flexible Neural Networks
This section explains the details of the proposed network, how information propagates through it, how synaptic strengths are updated, the architecture of the network, and how the property of structural flexibility is gained in the proposed method.
@@ 3.1. Information Propagation and Update of Synapses
The networks presented here consist of parameterized neural units and synaptic update rules. Except for input neurons, a neuron is represented by a small linear layer, and the hyperbolic tangent function as activation. Neural activations are thus small vectors rather than scalars. A synapse is represented by a vector of the same size as the neural activations. Propagation of activity from one neuron to another is done by element-wise multiplication of the activation of the pre-synaptic neuron and the synapse connecting it to the post-synaptic neuron. See Figure 1 for an illustration. Each synaptic vector is updated as being the hidden state of a GRU. The GRU takes the respective pre- and post-neurons’ activations as inputs and uses these to update its hidden state. The global reward given by the environment at the previous timestep is also concatenated with the input of the GRU. The synaptic GRUs of the network proposed here differ only from regular GRUs in that newly calculated hidden states are added to the previous hidden state instead of replacing it. A learning rate that is evolved alongside the gate parameters of the GRU is scaling the magnitude of the additions to the previous hidden state. The GRUs that make up the synapses of the network can be conceptualized as non-linear history-dependent plasticity rules, responsible for updating synaptic strengths.
@@
Figure 1. Pre- and post-neurons connected by synapse: Depicted is a representation of the flow of activity from one neuron to another through a synapse. Neural activations are small vectors, rather than just scalars. With the exception of the input neurons, each neuron is represented by a small linear layer with hyperbolic-tangent non-linearities. The output of the pre-neuron is modulated by the synapse via element-wise multiplication with the current values of the synaptic weights, and the resulting signal arrives as input to the post-neuron (solid black arrows). The output of both pre- and post-neurons plus the global reward signal is then used to compute the weight updates using the GRU (dotted arrows), which is then applied to the synaptic weights
3.2. Network Architecture
@@
Figure 2. Structurally Flexible Neural Network The input vector consists of the observation of the environment. This is sent through a reservoir layer with random connectivity. Neurons of a type are always pre-neurons to synapses of the same type. Each color corresponds to a distinct type that shares parameters. Each neuron has a linear layer associated with it, the parameters of which are shared by all other neurons of the same color. Likewise, each colored arrow corresponds to a synapse type. Each synapse type has one set of evolved GRU gate parameters associated with it. Importantly, even though the evolved gate parameters are shared between synapses of the same type, there is a unique hidden state associated with each single synapse, allowing for individual, history-dependent updates of each synapse.
@@
The adjacency matrix of the proposed networks is sparse, with half of the possible connections set to zero. For every new lifetime of an agent, a new adjacency matrix is sampled with each possible connection having an equal chance of being zero or not. As illustrated in Figure 2, connections can be formed from input neurons to the hidden neurons, between hidden neurons, from hidden neurons to output neurons, and from output neurons to hidden neurons. A connection that is set to zero at the beginning of the agent’s lifetime will not be updated. The initial synaptic weight values for each non-zero synapse are sampled from a uniform distribution, $\mathcal{U}$(-0.1, 0.1). Starting from this random initialization, the synapses be updated must through interactions with the environment such that a functional network emerges.
@@
At each timestep in the environment, all neurons of the network are updated multiple times through a number of micro ticks. The activation of each input neuron is set equal to an element of the observation from the environment. Since each neuron’s activation is a vector, the observation element of the respective input neuron is copied to each of the values of the neural activation. The information from the input neurons is sent through synaptic vectors to the hidden reservoir layer. The hidden neurons are updated with the signals from the input neurons as well as signals that were sent to them from hidden- or output neurons at the previous micro tick. Output neurons receive signals from hidden neurons at each micro tick. In all experiments below, two micro ticks were used for each environmental timestep. After the second micro tick, the activations of the output neurons are used to determine the action that the agent will carry out. The first elements of each of the output neurons’ activations are combined to form an action vector that is then passed to the environment. In the case of environments with discrete action spaces, the index of the largest value of this action vector is chosen as the action. At each micro tick, the first element of each of the signals from the output vectors to the hidden reservoir is set to correspond to the respective element of the action vector. If the action space of the environment is discrete, the first element of the signal from the output neuron with the index corresponding to the action taken is set to one and to zero for all others.
@@
Synapses are updated after the second micro tick. Activations from the previous environmental timestep are used as pre-activations to the synaptic GRUs, and the activations resulting from the second micro tick at the current environmental timestep are used as post-activations (see Figure 1).
@@ 3.3. Addressing The Symmetry Dilemma
The challenge that comes with constructing a parameterized approach that is symmetric and thus invariant to permutations and changes to the input or output size, is that the approach must be capable of resulting in information processing that is not symmetric, i.e., where the activation values that encode the flow of information through the symmetric architecture are themselves asymmetric.
With the network proposed in Section 3.1 and 3.2, we have the pieces to address the Symmetry Dilemma. The approach presented in this paper is greatly influenced by that of Kirsch and Schmidhuber  (Kirsch and Schmidhuber, 2020) and Kirsch et al.  (Kirsch et al., 2022a), but attempts to improve upon these through the introduction of parameterized neurons, and different sets of GRU parameters for each layer as described above.
As discussed by Pedersen and Risi  (Pedersen and Risi, 2022), the main requirement for permutation invariance is that no parameters can be optimized relative to a fixed position in the network. Along with parameter-sharing, this can avoided with the introduction of randomness in the network during evolution. Any aspect of the network structure that, if fixed throughout the evolution phase, would result in the evolved parameters overfitting to it must be presented with variation during the evolution phase for the parameters to be invariant to it. An example of such variation is that the hidden states of the synaptic GRUs can be initialized randomly to form an asymmetric structure in the network and still maintain permutation- and size-invariant properties. This is because the evolutionary optimization process has no information about these random initializations, and evolution does thus not rely on any specific initialization of the hidden states.
@@
As mentioned in Section 3.2, these two sources of randomness (random sparse connectivity and random weight initializations) are part of the approach proposed in this paper.
@@ 3.4. Neural and Synaptic Diversity
As noted by Tang and Ha  (Tang and Ha, 2021), all input units must share parameters to be invariant to permutations of the input elements. However, synapses from the input to a hidden layer, do not need to share parameters with synapses between other layers to adhere to the requirements for invariance. With the proposed SFNN approach of this paper, only synapses that connect the same inputs to the same outputs must share parameters.
@@ 4. Experiments
In this section, we describe experiments to test the usefulness of having distinct parameterizations of synapses and neurons of different layers that are sparsely connected. Since the main point of SFNNs is that they are not limited to a specific input- and output space, we optimize the SFNN parameters on three different environments at the same time.
@@ 4.1. Environments
We test our model in multiple simple control tasks (Fig. 3) implemented in the Gymnax suite (Lange, 2022).
In all experiments, the ordering of the inputs to and outputs from the models stays fixed throughout the entire optimization time.
@@ 4.2. Training Setup
The training of the type of network described in Section 3 is set up as described in the following.
Similar to the setup of Kirsch et al.  (Kirsch et al., 2022a), and Bertens and Lee  (Bertens and Lee, 2020), the fitness of an individual is determined over a lifetime that spans multiple episodes in a given environment. However, since we are evolving parameters on three tasks, an individual actually goes through three lifetimes, one for each environment. Beyond the differences in inputs, the agent receives no information about which environment it is placed in at the beginning of a new lifetime.
@@
For each episode during the lifetime of the agent, the accumulated reward for each time step is recorded. The lifetime score of the agent is then the weighted sum of recorded episode scores. Importantly, the state of the network of the agent is carried over from episode to episode, so the agent has the chance to learn throughout its lifetime.
Scores obtained in episodes later in the agent’s lifetime are weighted higher than scores obtained early in life. Specifically, the score of each episode is weighted by the enumeration of the episode divided by the sum of all episode enumerations. The reason for assigning a larger weight to later episodes is to encourage learning over time and to maintain performance. Agents that perform well in the first half of their lifetimes, but then start to fail, should be differentiated from agents that with experience are able to improve their scores. Letting scores from the end of the agent’s lifetime be more important for the lifetime score also serves to not punish early exploration.
@@
The parameters of the neural and synaptic units are optimized using CMA-ES (Hansen, 2006, 2016). In all experiments, the population size is set to 128.
The specifics of the SFFN approach of the main experiments are summarized in Table 1. The linear layer of a neuron is in this case a four-by-four weight matrix. The hidden size of the GRUs is four as well, and the input to the GRU is thus two times four plus the one global reward element. For each set of GRU parameters, a single scalar is also optimized to be used as a learning rate for the hidden state updates. In practice, the learning rate is always multiplied by a constant of 0.01 to form an applied learning rate. This is done to ensure that the magnitude of the applied learning rate is smaller than the rest of the evolved parameters, at least at the beginning of the evolution phase.
@@
Figure 4 provides an overview of how many evolved and plastic parameters each setting has. For all settings, five evolution runs with different seeds are carried out.
Figure 4. Number of evolved and plastic parameters: The different trained versions differ in their numbers of evolved and plastic parameters. Evolved parameters are the synaptic GRU parameters and the linear layers of the neurons. The plastic parameters are the synaptic weight values that are updated by the GRUs (LSTMs in the case of SymLA). The SFNN_single version has the same number of plastic parameters as SFNN, but fewer evolved ones, as only parameters for a single neuron and synapse type are evolved. The SFNN_fully version has the same number of parameters as SFNN, but has double the amount of plastic parameters, as no synapses are excluded. The SymLA method has many more evolved parameters than plastic, as the synaptic units are much bigger than in SFNN and there are fewer of them, as the SymLA networks are shallow.

### v2

@@ Abstract.
Artificial neural networks used for reinforcement learning are structurally rigid, meaning that each optimized parameter of the network is tied to its specific placement in the network structure. Structural rigidity limits the ability to optimize parameters of policies across multiple environments that do not share input and output spaces. This is a consequence of the number of optimized parameters being directly dependent on the structure of the network. In this paper, we present Structurally Flexible Neural Networks (SFNNs), which consist of connected gated recurrent units (GRUs) as synaptic plasticity rules and linear layers as neurons. In contrast to earlier work, SFNNs contain several different sets of parameterized building blocks. Here we show that SFNNs can overcome the challenging symmetry dilemma, which refers to the problem of optimizing units with shared parameters to each express different representations during deployment. In this paper, the same SFNN can learn to solve three classic control environments that have different input/output spaces. SFFNs thus represent a step toward a more general model capable of solving several environments at once.
@@ 1. Introduction
The introduction of deep neural networks has led to remarkable improvements when it comes to the control of artificial agents either through optimization using reinforcement learning (Mnih et al., 2013; Arulkumaran et al., 2017) or evolutionary algorithms (Salimans et al., 2017; Stanley et al., 2019). One key finding has been that more diversity in the training environments leads to policies with stronger generalization capabilities (e.g., (Won and Lee, 2019; Kumar et al., 2021; Kirsch et al., 2022b; Risi and Togelius, 2020)). In the ideal case, we would optimize our agents across a wide array of training environments with differing features, including different input and output spaces. However, most neural networks are structurally rigid. A consequence of this inflexibility is that their architectures are tied to the specifics of a particular input and output space, which prevents typical neural networks from being optimized across domains with differing dimensions.
A key challenge on the road to achieving this goal is overcoming what we refer to as the Symmetry Dilemma. The Symmetry Dilemma is the tension between, on the one hand, gaining desirable properties that come from inducing symmetry in the optimized parameters, and on the other hand, maintaining a structure that is capable of distributing information across its nodes. If we optimize parameters to form a symmetric structure, meaning that any parameter could be placed anywhere within a network structure without changing the functionality of the network, we gain structurally flexible properties such as invariance to permutations and changes in the sizes of the input and/or output (Tang and Ha, 2021; Kirsch et al., 2022a; Kirsch et al., 2022b). However, to distribute information within the structure, each parameter must be able to contain or express individualized information.
@@
This paper presents a novel approach called Structurally Flexible Neural Networks (SFNN). The main features of these networks are that they consist of sparsely connected parameterized neurons and dynamic synapses. The optimized parameters are thus those of the building blocks of a neural network, separated from the network structure. Special for SFNNs, the network contains several different sets of parameterized building blocks. These are referred to as different types of neurons and synapses.
The experiments in this paper showcase a single parameter set capable of solving tasks with different input and output sizes. Starting with random connectivity, and random initial synaptic weight values, the evolved plastic synapses and neurons organize during the lifetime to achieve better performance in the given environment. The key to achieving a versatile learning algorithm lies in structural flexibility, which liberates the model from being tied to specific input and output spaces. The ultimate objective is to devise a set of building blocks, that when put together to form a network, exhibit fast adaptation across various environments, regardless of their dimensionality, and the permutations of the input and output elements. Recently, the notion of foundation models has surfaced in the context of reinforcement learning (Yang et al., 2023). A structurally flexible model could serve as a general foundation model (Bommasani et al., 2021; Yang et al., 2023) for reinforcement learning (RL)-type tasks, even when the tasks do not share input and output dimensions.
@@ 2. Related Work
Plasticity as Indirect Encoding. Our approach uses Gated Recurrent Units (GRUs) (Chung et al., 2014) with shared parameters to update synaptic weight values. As such, it is related to approaches that optimize plasticity mechanisms rather than directly optimizing weights. Indirect encodings offer a powerful way to compress the specification of large systems - like the weights of a neural network - using smaller functions (Zador, 2019; Pedersen and Risi, 2021; Stanley and Miikkulainen, 2003). From this perspective, approaches that optimize plasticity mechanisms rather than directly optimizing weights can be seen as a way to encode multiple realizations of the parameters of a neural network since these can take different values at different points during the lifetime of an agent (Najarro and Risi, 2020).
@@
Graph Neural Networks. The network type presented in this paper consists of synapses that are themselves recurrent neural networks. This means that these synapses all have hidden state vectors associated with them, and, as explained in Section 3.1, these are updated using local information as well as a global reward signal. In this regard, they share similarities in how Graph Neural Networks (GNNs) work. GNNs are designed to analyze data represented as graphs (Wu et al., 2020). Nodes in a GNN have internal states and pass information to their neighbors through parameterized message functions that are optimized during the network’s training phase (Zhou et al., 2020). A neural network with parameterized neurons and synapses being GRUs can be seen as a graph where some nodes (the synapses) can only have two neighbors (neurons), whereas the ”neuron” nodes can have many ”synaptic” nodes as neighbors. In the networks in this paper, these different types of nodes also differ in how they propagate information and update their states. However, the idea of optimizing how nodes with internal states interact with each other is analogous to that of GNNs. GNNs have been used to control simulated robots showing that they can take advantage of the fact that robot morphology can be expressed as a graph (Wang et al., 2018). In contrast, the networks presented in this paper do not make assumptions about the spatial structure of inputs or outputs of the model but are only concerned with how the units of the network can best be updated.
@@ 3. Structurally Flexible Neural Networks
This section explains the details of the proposed structural flexible neural network approach, including how the information propagates works, how synaptic strengths are updated, the architectural details, and how the property of structural flexibility is gained.
@@ 3.1. Information Propagation and Update of Synapses
The networks consist of parameterized neural units and synaptic update rules. A neuron is represented by a small linear layer, and the hyperbolic tangent as activation function. Neural activations are thus small vectors rather than scalars. A synapse is represented by a vector of the same size as the neural activations. Propagation of activity from one neuron to another is done by element-wise multiplication of the activation of the pre-synaptic neuron and the synapse connecting it to the post-synaptic neuron (Figure 1). Each synaptic vector is updated as being the hidden state of a GRU. The GRU takes the respective pre- and post-neurons’ activations as inputs and uses these to update its hidden state. The global reward given by the environment at the previous timestep is also concatenated with the input of the GRU. The synaptic GRUs of the network proposed here differ only from regular GRUs in that newly calculated hidden states are added to the previous hidden state instead of replacing it. A learning rate that is evolved alongside the gate parameters of the GRU is scaling the magnitude of the additions to the previous hidden state. The GRUs that make up the synapses of the network can be conceptualized as non-linear history-dependent plasticity rules, responsible for updating synaptic strengths.
@@
Figure 1. (A) Building Blocks of Structurally Flexible Neural Networks Depicted is a representation of the flow of activity from one neuron to another through a synapse. Neural activations are small vectors, rather than just scalars. Each neuron is represented by a small linear layer with hyperbolic-tangent non-linearities, similar to (Pedersen and Risi, 2023). The output of the pre-neuron is modulated by the synapse via element-wise multiplication with the current values of the synaptic weights, and the resulting signal arrives as input to the post-neuron (solid black arrows). The output of both pre- and post-neurons plus the global reward signal is then used to compute the weight updates using the GRU (dotted arrows), which is then applied to the synaptic weights. (B) Structurally Flexible Neural Network. The input vector consists of the observation from the environment, which is sent through a reservoir layer with random connectivity. Neurons of a certain type are always pre-neurons to synapses of the same type. Each color corresponds to a distinct type (input, hidden, output) that shares parameters. Each neuron has a linear layer associated with it, the parameters of which are shared by all other neurons of the same color. Likewise, each colored arrow corresponds to a synapse type. Each synapse type has one set of evolved GRU gate parameters associated with it. Importantly, even though the evolved gate parameters are shared between synapses of the same type, there is a unique hidden state associated with each single synapse, allowing for individual, history-dependent updates of each synapse.
@@
The adjacency matrix of the proposed networks is sparse, with half of the possible connections set to zero. The proportion of connections set to zero is a hyperparameter that could take other values or be optimized. For every new lifetime of an agent, a new adjacency matrix is sampled with each possible connection having an equal chance of being zero or not. As illustrated in Figure 1, connections can be formed from input neurons to the hidden neurons, between hidden neurons, from hidden neurons to output neurons, and from output neurons to hidden neurons. A connection that is set to zero at the beginning of the agent’s lifetime will not be updated. The initial synaptic weight values for each non-zero synapse are sampled from a uniform distribution, $\mathcal{U}$(-0.1, 0.1). Starting from this random initialization, the synapses must be updated through interactions with the environment such that a functional network emerges.
@@
At each timestep in the environment, all neurons of the network are updated multiple times through a number of micro ticks. The activation of each input neuron is set equal to an element of the observation from the environment. Since each neuron’s activation is a vector, the observation element of the respective input neuron is copied to each of the values of the neural activation. The information from the input neurons is sent through synaptic vectors to the hidden reservoir layer. The hidden neurons are updated with the signals from the input neurons as well as signals that were sent to them from hidden- or output neurons at the previous micro tick.
Output neurons receive signals from hidden neurons at each micro tick. In all experiments below, two micro ticks were used for each environmental timestep. After the second micro tick, the activations of the output neurons are used to determine the action that the agent will carry out. The first elements of each of the output neurons’ activations are combined to form an action vector that is then passed to the environment. For environments with discrete action spaces, the index of the largest value of this action vector is chosen as the action. At each micro tick, the first element of each of the signals from the output vectors to the hidden reservoir is set to correspond to the respective element of the action vector. The first element of the signal from the output neuron with the index corresponding to the discrete action taken is set to one and to zero for all others.
@@
Synapses are updated after the second micro tick. Activations from the previous environmental timestep are used as pre-activations to the synaptic GRUs, and the activations resulting from the second micro tick at the current environmental timestep are used as post-activations.
@@ 3.2. Addressing The Symmetry Dilemma
Symmetric parameterization means that it is possible to switch around the placement of parameters without changing the function. The challenge that comes with constructing a parameterized approach that is symmetric and thus invariant to permutations and changes to the input or output size, is that the approach must be capable of resulting in information processing that is not symmetric, i.e., where the activation values that encode the flow of information through the symmetric architecture are themselves asymmetric.
With the network proposed in Section 3.1 and 1, we have the pieces to address the Symmetry Dilemma. The approach presented in this paper is greatly influenced by that of Kirsch and Schmidhuber  (Kirsch and Schmidhuber, 2020) and Kirsch et al.  (Kirsch et al., 2022a), but attempts to improve upon these through the introduction of parameterized neurons, and different sets of GRU parameters for each layer as described above.
As discussed by Pedersen and Risi  (Pedersen and Risi, 2022), the main requirement for permutation invariance is that no parameters can be optimized relative to a fixed position in the network. Along with parameter-sharing, this can be avoided with the introduction of randomness in the network during evolution. Any aspect of the network structure that, if fixed throughout evolution, would result in the evolved parameters overfitting to it must be presented with variation during the evolution phase for the parameters to be invariant to it.
An example of such variation is that the hidden states of the synaptic GRUs can be initialized randomly to form an asymmetric structure in the network and still maintain permutation- and size-invariant properties. This is because the evolutionary optimization process has no information about these random initializations, and evolution does thus not rely on any specific initialization of the hidden states.
@@
As mentioned in Section 1, these two sources of randomness (random sparse connectivity and random weight initializations) are part of the approach proposed in this paper.
@@ 3.3. Neural and Synaptic Diversity
As noted by Tang and Ha  (Tang and Ha, 2021), all input units must share parameters to be invariant to permutations of the input elements. However, synapses from the input to a hidden layer do not need to share parameters with synapses between other layers to adhere to the requirements for invariance. With the proposed SFNN approach of this paper, only synapses that connect the same inputs to the same outputs must share parameters.
@@ 4. Experiments
In this section, we describe the experiments to test the usefulness of having distinct parameterizations of synapses and neurons of different layers that are sparsely connected. Since the main point of SFNNs is that they are not limited to a specific input- and output space, we optimize the SFNN parameters on three different environments at the same time.
@@ 4.1. Environments
We test our model in multiple simple control tasks (Figure 2) implemented in the Gymnax suite (Lange, 2022). In all experiments, the ordering of the inputs to and outputs from the models stays fixed throughout the entire optimization time.
@@ 4.2. Training Setup
This section details the training setup for SFNNs in this paper. Similar to the setup of Kirsch et al.  (Kirsch et al., 2022a), and Bertens and Lee  (Bertens and Lee, 2020), the fitness of an individual is determined over a lifetime that spans multiple episodes in a given environment. However, since we are evolving parameters on three tasks, an individual actually goes through three lifetimes, one for each environment. Beyond the differences in inputs, the agent receives no information about which environment it is placed in at the beginning of a new lifetime.
@@
For each episode during the agent’s lifetime, the accumulated reward for each time step is recorded. The lifetime score of the agent is then the weighted sum of recorded episode scores. Importantly, the state of the network of the agent is carried over from episode to episode, so the agent has the chance to learn throughout its lifetime. Scores obtained in episodes later in the agent’s lifetime are weighted higher than ones obtained early in life. Specifically, the score of each episode is weighted by the enumeration of the episode divided by the sum of all episode enumerations. The reason for assigning a larger weight to later episodes is to encourage learning over time and to maintain performance. Agents that perform well in the first half of their lifetimes, but then start to fail, should be differentiated from agents who improve their scores with experience. Letting scores from the end of the agent’s lifetime be more important for the lifetime score also serves to not punish early exploration.
@@
The parameters of the neural and synaptic units are optimized using CMA-ES (Hansen, 2006, 2016). In all experiments, the population size is set to 128. The specifics of the SFFN approach of the main experiments are summarized in Table 1. The linear layer of a neuron is in this case a four-by-four weight matrix. The hidden size of the GRUs is four as well, and the input to the GRU is thus two times four plus the one global reward element. For each set of GRU parameters, a single scalar is also optimized to be used as a learning rate for the hidden state updates. In practice, the learning rate is always multiplied by a constant of 0.01 to form an applied learning rate. This is done to ensure that the magnitude of the applied learning rate is smaller than the rest of the evolved parameters, at least at the beginning of the evolution phase.
@@
Figure 3 provides an overview of how many evolved and plastic parameters each setting has. For all settings, five evolutionary runs with different seeds were carried out.
Figure 3. Number of evolved and plastic parameters: The approaches differ in their numbers of evolved and plastic parameters. Evolved parameters are the synaptic GRU parameters and the linear layers of the neurons. The plastic parameters are the synaptic weight values that are updated by the GRUs (LSTMs in the case of SymLA). The SFNN_single version has the same number of plastic parameters as SFNN, but fewer evolved ones, as only parameters for a single neuron and synapse type are evolved. The SFNN_fully version has the same number of parameters as SFNN, but has double the amount of plastic parameters, as no synapses are excluded. The SymLA method has many more evolved parameters than plastic, as the synaptic units are much bigger than in SFNN and there are fewer of them, as the SymLA networks are shallow.